# Optimal Procurement Design: A Reduced Form Approach

### v1

@@ 4.1 Optimal interim allocation
$\mathcal{L}_{\gamma}=\int_{0}^{1}\left[\widetilde{h}_{\gamma}(s)+\lambda\widetilde{g}(s)\right]\widetilde{P}(s)\,\mathrm{d}s,$
where $\widetilde{h}_{\gamma}(s):=h_{\gamma}(F^{-1}(s))$ is the quantile version of the weighted virtual surplus. Define $\phi_{\gamma}(q;\lambda):=h_{\gamma}(q)+\lambda g(q)$, and let $\widetilde{\phi}_{\gamma}(s;\lambda):=\phi_{\gamma}(F^{-1}(s);\lambda)$. Evidently, $\widetilde{\phi}_{\gamma}(s;\lambda)=\widetilde{h}_{\gamma}(s)+\lambda\widetilde{g}(s)$, which is the quantile virtual surplus of the Lagrangian.
We solve the problem by maximizing the Lagrangian over all $\widetilde{P}\in\Omega_{w}(\widetilde{P}^{*})$, and then find an appropriate Lagrangian multiplier such that the complimentary slackness condition holds. More specifically, we iron $\phi_{\gamma}(s;\lambda)$ to make sure that the monotonicity constraint holds: let
$\Phi_{\gamma}(s;\lambda):=\int_{0}^{s}\widetilde{\phi}_{\gamma}(s;\lambda)\,\mathrm{d}x,$
and
$\overline{\phi}_{\gamma}(s;\lambda):=\frac{\partial}{\partial s}\,\overline{\Phi}_{\gamma}(s;\lambda)$
is the ironed quantile virtual surplus of the Lagrangian. Similar to Proposition 1, for all $s$ such that $\overline{\phi}_{\gamma}(s;\lambda)>0$, the optimal interim allocation $\hat{P}$ is flat whenever ironing is needed, and coincide with $\widetilde{P}^{*}(s)=(1-s)^{n-1}$ otherwise. If there exists an interval $[S_{+},S_{0}]$ on which $\overline{\phi}_{\gamma}(s;\lambda)=0$, we may need to find some $\overline{P}$ satisfying
@@
and set $\hat{P}(s)=\overline{P}$ on $[S_{+},S_{0}]$ to satisfy complementary slackness; the second inequality in the above expression is needed because Border’s condition requires $\hat{P}\prec_{w}\widetilde{P}^{*}$.
@@
If there exist $\lambda^{*}\geq 0$, $0\leq S_{+}\leq S_{0}\leq 1$, and a collection of disjoint intervals $\{\left[\underline{s}_{i},\bar{s}_{i}\right)\}_{i\in\mathcal{I}}$ with $\left[\underline{s}_{i},\bar{s}_{i}\right)\subseteq[0,S_{0}]$ for each $i\in\mathcal{I}$, such that
@@
    $S_{+}=\sup\left\{s\in[0,1]:\overline{\phi}_{\gamma}(s;\lambda^{*})>0\right\}$ and $S_{0}=\sup\left\{s\in[0,1]:\overline{\phi}_{\gamma}(s;\lambda^{*})\geq 0\right\}$;
@@
    $\overline{\Phi}_{\gamma}(s;\lambda^{*})$ is affine on $\left[\underline{s}_{i},\bar{s}_{i}\right)$ for each $i\in\mathcal{I}$ and $[S_{+},S_{0})$, and
@@
    $\overline{\Phi}_{\gamma}(s;\lambda^{*})=\Phi_{\gamma}(s;\lambda^{*})$ on $[0,S_{+}]\big{/}\bigcup_{i\in\mathcal{I}}\left[\underline{s}_{i},\overline{s}_{i}\right)$,
@@
$\hat{P}(s)=\left\{\begin{array}[]{ll}(1-s)^{n-1}&\text{if }s\in[0,S_{+}]\big{/}\bigcup_{i\in\mathcal{I}}\left[\underline{s}_{i},\overline{s}_{i}\right)\\ \frac{\int_{\underline{s}_{i}}^{\overline{s}_{i}}(1-s)^{n-1}\mathrm{d}s}{\overline{s}_{i}-\underline{s}_{i}}&\text{if }s\in\left[\underline{s}_{i},\overline{s}_{i}\right)\\ \overline{P}&\text{if }s\in(S_{+},S_{0}]\\ 0&\text{if }s\in(S_{0},1]\end{array}\right.$ (11)
with $\overline{P}\in\left(0,A(S_{+},S_{0})\right]$ satisfying the complementary slackness condition
@@
4.2 Implementation: Augmenting BRA if necessary
If the optimal interim allocation $\hat{P}$ defined in (11) satisfies $S_{+}=S_{0}$, it takes exactly the same form as (9) in Proposition 1. Therefore, its implementation is described by Theorem 1, by replacing $\overline{S}$ by $S_{0}$. If, instead, $S_{+}<S_{0}$, then to implement the optimal interim allocation we need to slightly modify the BRA. In this modified version, if a seller with quality $q\in(S_{+},S_{0}]$ wins the auction, she supplies the good with probability $\overline{P}/A(S_{+},S_{0})$. The details of this modification are described below.
@@ 4.2 Implementation: Augmenting BRA if necessary
An augmented bid-restricted auction (aBRA) is a sealed bid auction with
@@
  • With probability $1-\zeta$, sellers who bid $B$ are disqualified, meaning that their bids do not count.
@@
  • The seller whose bid is the lowest wins the auction; in the event of a tie, the winning seller is chosen uniformly at random.
@@
  • If the winning bid is the only bid in its bid interval $[\underline{b}_{i},\overline{b}_{i}]$ for some $i=1,\ldots,M$, and
@@
    • if, furthermore, the second-lowest bid equals $\underline{b}_{j}$ for some $j>i$ with $k$ other sellers bidding $\underline{b}_{j}$, then the winning seller receives a payment of $(\underline{b}_{j}+k\overline{b}_{j-1})/(k+1)$;

### v2

@@ C.1 Optimal interim allocation
$\mathcal{L}_{\gamma}(\widetilde{P};\lambda)\coloneqq\int_{0}^{1}\left[\widetilde{h}_{\gamma}(s)+\lambda\widetilde{g}(s)\right]\widetilde{P}(s)\,\mathrm{d}s,$
[…]
We iron $\widetilde{\phi}_{\gamma}(s;\lambda)$ to make sure that the monotonicity constraint holds: let
$H_{\gamma}(s)\coloneqq\int_{0}^{s}\widetilde{h}_{\gamma}(x)\,\mathrm{d}x\quad\text{and}\quad\Phi_{\gamma}(s;\lambda)\coloneqq\int_{0}^{s}\widetilde{\phi}_{\gamma}(x;\lambda)\,\mathrm{d}x,$
and let $\overline{H}_{\gamma}$ and $\overline{\Phi}_{\gamma}$ be the concave hulls of $H_{\gamma}$ and $\Phi_{\gamma}$, respectively. Now let
$\overline{h}_{\gamma}(s)\coloneqq\overline{H}_{\gamma}^{\prime}(s),\quad\text{and}\quad\overline{\phi}_{\gamma}(s;\lambda)\coloneqq\frac{\partial}{\partial s}\,\overline{\Phi}_{\gamma}(s;\lambda);$
$\overline{h}_{\gamma}(s)$ and $\overline{\phi}_{\gamma}(s;\lambda)$ are the ironed quantile weighted virtual surplus and the ironed quantile virtual surplus of the Lagrangian, respectively. By construction, both are decreasing in $s$.
Similar to the buyer-optimal case in Proposition 1, for all $s$ such that $\overline{\phi}_{\gamma}(s;\lambda)>0$, the optimal interim allocation $\hat{P}$ is flat whenever ironing is needed, and coincides with $\widetilde{P}^{*}(s)=(1-s)^{n-1}$ otherwise. If there exists an interval $S_{+},S_{0}]$ on which $\overline{\phi}_{\gamma}(s;\lambda)=0$, we may need to find some $\overline{P}$ satisfying
@@
and set $\hat{P}(s)=\overline{P}$ on $S_{+},S_{0}]$ to satisfy complementary slackness; the second inequality in the above expression is needed because Border’s condition requires $\hat{P}\prec_{w}\widetilde{P}^{*}$.
@@
If there exist $\lambda^{*}\geq 0$, $0\leq S_{+}\leq S_{0}\leq 1$, and a collection of disjoint intervals $\{\left[\underline{s}_{i},\overline{s}_{i}\right)\}_{i\in\mathcal{I}}$ with $\left[\underline{s}_{i},\overline{s}_{i}\right)\subseteq[0,S_{0}]$ for each $i\in\mathcal{I}$, such that
@@
    $S_{+}=\sup\left\{s\in[0,1]:\overline{\phi}_{\gamma}(s;\lambda^{*})>0\right\}$ and $S_{0}=\sup\left\{s\in[0,1]:\overline{\phi}_{\gamma}(s;\lambda^{*})\geq 0\right\}$,
@@
    $\overline{\Phi}_{\gamma}(s;\lambda^{*})$ is affine on $\left[\underline{s}_{i},\overline{s}_{i}\right)$ for each $i\in\mathcal{I}$ and $S_{+},S_{0})$, and
@@
[…]
@@
$\hat{P}(s)=\left\{\begin{array}[]{ll}(1-s)^{n-1}&\text{if }s\in[0,S_{+}]\big\backslash\bigcup_{i\in\mathcal{I}}\left[\underline{s}_{i},\overline{s}_{i}\right)\\ \frac{\int_{\underline{s}_{i}}^{\overline{s}_{i}}(1-s)^{n-1}\mathrm{d}s}{\overline{s}_{i}-\underline{s}_{i}}&\text{if }s\in\left[\underline{s}_{i},\overline{s}_{i}\right)\\ \overline{P}&\text{if }s\in(S_{+},S_{0}]\\ 0&\text{if }s\in(S_{0},1]\end{array}\right$. (11)
with $\overline{P}\in\left(0,A(S_{+},S_{0})\right]$ such that $\hat{P}$ satisfies the complementary slackness condition
@@
Proof.
Suppose first that $\overline{h}_{\gamma}(0)<0$, meaning that it is undesirable to trade under incomplete information. In this case, the BRA with one (trivial) bid interval $\{0\}$ is optimal. Now suppose instead that $\overline{h}_{\gamma}(0)\geq 0$. Define $\left\{s\in[0,1]:\overline{h}_{\gamma}(s)=0\right\}; ZZ$ is the set of points on which the ironed quantile social surplus is zero. Because $\overline{h}_{\gamma}(0)\geq 0$, if $Z$ is empty, we must have $\overline{h}_{\gamma}(1)>0$.
[…]
$\lambda\int_{0}^{1}\widetilde{g}(s)\hat{P}(s)\,\mathrm{d}s=0$
Step 1.
Check whether there exists $\hat{s}\in[0,1]$, either $\hat{s}\in Z$, or $Z=\varnothing$ and $\hat{s}=1$ such that
$\hat{P}(s)=\left\{\begin{array}[]{ll}(1-s)^{n-1}&\text{if }s\in[0,\hat{s}]\big\backslash\bigcup_{i\in I}\left[\underline{x}_{i},\overline{x}_{i}\right)\\ \frac{\int_{\underline{x}_{i}}^{\overline{x}_{i}}(1-s)^{n-1}\mathrm{d}s}{\overline{x}_{i}-\underline{x}_{i}}&\text{if }s\in\left[\underline{x}_{i},\overline{x}_{i}\right)\\ 0&\text{if }s\in(\hat{s},1]\end{array}\right$.
satisfies
$\int_{0}^{1}\widetilde{g}(s)\hat{P}(s)\,\mathrm{d}s\geq 0,$
where $\left\{\left[\underline{x}_{i},\overline{x}_{i}\right)\right\}_{i\in I}$ is the collection of intervals on which $\overline{H}_{\gamma}$ is affine. If so, we can set $\lambda=0$, which implies that the ironed quantile virtual surplus of the Lagrangian coincides with the ironed quantile weighted virtual surplus: $\overline{\phi}_{\gamma}(s;0)=\overline{h}_{\gamma}(s)$. Setting $S_{+}=S_{0}=\hat{s}$, $\hat{P}$ solves problem (5). If not, go to Step 2.
Step 2.
We must have $\lambda>0$, otherwise we could have found an $\hat{s}$ in Step 1. Now we search for $\lambda>0$ such that there exists a unique $\tilde{s}$ such that $\overline{\phi}_{\gamma}(\tilde{s};\lambda)=0$, and the “induced interim allocation”
$\hat{P}(s)=\left\{\begin{array}[]{ll}(1-s)^{n-1}&\text{if }s\in[0,\tilde{s}]\big\backslash\bigcup_{i\in J}[\underline{y}_{i},\overline{y}_{i})\\ \frac{\int_{\underline{y}_{i}}^{\overline{y}_{i}}(1-s)^{n-1}\mathrm{d}s}{\overline{y}_{i}-\underline{y}_{i}}&\text{if }s\in[\underline{y}_{i},\overline{y}_{i})\\ 0&\text{if }s\in(\tilde{s},1]\end{array}\right$.
satisfies
$\int_{0}^{1}\widetilde{g}(s)\hat{P}(s)\,\mathrm{d}s=0,$
[…]
Step 3.
There must exist an interval $S_{+},S_{0}]\subseteq[0,1]$ with $S_{+}<S_{0}$ such that $\overline{\phi}_{\gamma}(\,\cdot\,;\lambda)=0$ on $S_{+},S_{0}]$, and there exists $\overline{P}$ with
such that
$\hat{P}(s)=\left\{\begin{array}[]{ll}(1-s)^{n-1}&\text{if }s\in[0,S_{+}]\big\backslash\bigcup_{i\in J}[\underline{y}_{i},\overline{y}_{i})\\ \frac{\int_{\underline{y}_{i}}^{\overline{y}_{i}}(1-s)^{n-1}\mathrm{d}s}{\overline{y}_{i}-\underline{y}_{i}}&\text{if }s\in[\underline{y}_{i},\overline{y}_{i})\\ \overline{P}&\text{if }s\in(S_{+},S_{0}]\\ 0&\text{if }s\in(S_{0},1]\end{array}\right$.
satisfying
$\int_{0}^{1}\widetilde{g}(s)\hat{P}(s)\,\mathrm{d}s=0$
solves problem (5). ∎
C.2 Implementation: Augmenting BRA if necessary
If the optimal interim allocation $\hat{P}$ defined in (11) satisfies $S_{+}=S_{0}$, it takes exactly the same form as (8) in Proposition 1. Therefore, its implementation is described by Theorem 1, by replacing $\overline{S}$ by $S_{0}$. If, instead, $S_{+}<S_{0}$, then to implement the optimal interim allocation we need to slightly modify the BRA. The details of this modification are described below.
@@ C.2 Implementation: Augmenting BRA if necessary
An augmented bid-restricted auction (aBRA) is a sealed-bid auction with
@@
  • With probability $1-\zeta$, all sellers who bid $B$ are disqualified, meaning that their bids do not count.
@@
  • The seller whose (valid) bid is the lowest wins the auction; in the event of a tie, the winning seller is chosen with equal probability among the sellers submitting the lowest (valid) bid.
@@
  • If the winning bid is the only bid in its standard bid interval $\underline{b}_{i},\overline{b}_{i}]$ for some $i=1,\ldots,M$, and
@@
    • if, furthermore, the second-lowest (valid) bid equals $\underline{b}_{j}$ for some $j>i$ and a total of $k$ sellers bid $\underline{b}_{j}$, then the winning seller receives a payment of $(\underline{b}_{j}+k\overline{b}_{j-1})/(k+1)$;